High-purity benzoic acid (C6H5COOH; ΔHrxn for combustion = −3227 kJ/mol) is used as a standard for calibrating bomb calorimeters. A 1.221-g sample burns in a calorimeter (heat capacity = 1365 J/°C) that contains exactly 1.370 kg of water. What temperature change is observed?

1
Expert's answer
2021-03-08T06:11:14-0500

Sample of benzoic acid =1.221 g,and C6H5COOH; ΔHrxn for combustion = −3227 kJ/mol) and calorimeter heat capacity = 1365 J/°C which has exactly 1.37 Kg of water

Firstly we calculate mole of benzoic acid

number of moles, "n= \frac{given mass}{molar mass}= \frac{1.221}{122.1}=0.01"

Now enthalpy of combustion="(0.01 \times 3227)=32.27 KJ"

As per principle of calorimetry

Heat released by combustion= heat absorb by calorimeter+ heat absorb by water

"32270 = q_1 \Delta T + q_2 \Delta T" (q2= mc) here c = specific heat of water= 4.2 J

"32270 = 1365 \Delta T + 1370 \times 4.2 \Delta T"

"32270 = 7119 \Delta T"

"\Delta T=4.533 K" This is change in temperature